what is a solution to 8cosx+4=0 in degrees?
a. undefined
b.120*
c. -30*
d. 1*

Answer by jim_thompson5910(35256)

8cos(x)+4=0

8cos(x)=-4

cos(x)=-4/8

cos(x)=-1/2

I'll let you take over from here.